# Generative Evolutionary Meta-Solver (GEMS): Scalable Surrogate-Free Multi-Agent Learning

### v1

@@ 3.5 Training the Generator with Amortized Best Response
Once a promising latent code $z_{t}^{\star}$ has been identified, Gems must ensure that the generator $G_{\theta}$ can realize the associated high-performing policy while retaining its ability to produce previously effective policies. This is achieved via an Amortized Best-Response with a Trust Region (ABR-TR) objective, inspired by trust-region methods in deep reinforcement learning such as TRPO and PPO ([16, 17]).
Rather than training a new network from scratch, Gems fine-tunes the existing generator $G_{\theta}$ to maximize the performance of a curated set of promising latent codes against the opponent mixture $\sigma_{t+1}$. A KL-divergence penalty against a frozen, older generator $\theta^{-}$ serves as a trust region, mitigating catastrophic forgetting:
$\mathcal{L}_{\mathrm{ABR-TR}}(\theta)\;=\;\mathbb{E}_{z\sim q_{t},\,j\sim\sigma_{t+1}}\!\Bigl[\widehat{A}_{\pi_{G_{\theta}(z)}}\;-\;\beta\,D_{\mathrm{KL}}\!\bigl(\pi_{G_{\theta}(z)}\;\|\;\pi_{G_{\theta^{-}}}(z)\bigr)\;-\;\lambda_{J}\|JG_{\theta}(z)\|_{F}^{2}\Bigr],$ (8)
where $\widehat{A}$ denotes a suitable advantage estimator. Taking a few gradient-ascent steps on $\mathcal{L}_{\mathrm{ABR-TR}}$ each iteration amortizes best-response computation while preserving a stable, single generator.
@@ 3.6 Overall Exploitability Bound
The preceding components now combine to yield a convergence guarantee for the entire Gems algorithm. The average exploitability decomposes into four interpretable terms, each tracing back to a distinct error source: (1) the inherent regret of the OMWU meta-strategy solver; (2) noise from Monte-Carlo estimation; (3) sub-optimality of the bandit oracle; and (4) approximation error introduced by amortized generator training.
Let $\varepsilon_{\mathrm{BR}}$ denote the gap between a true best response and the policy produced by the generator for the selected latent code $z_{t}^{\star}$:
$\varepsilon_{\mathrm{BR}}\;=\;\sup_{t}\mathbb{E}_{j\sim\sigma_{t}}\!\Bigl[r\bigl(\pi_{\mathrm{BR}}(j),\pi_{j}\bigr)\;-\;r\bigl(\pi_{G_{\theta}(z_{t}^{\star})},\pi_{j}\bigr)\Bigr]\;\geq 0.$ (9)
@@
Assume rewards lie in $[0,1]$ and that oracle selection operates on a two-time-scale schedule. With OMWU step size $\eta=\Theta\!\bigl(\sqrt{\ln k_{T}/T}\bigr)$ and a Monte-Carlo budget satisfying $\mathbb{E}[\|\hat{v}_{t}-M\sigma_{t}\|_{\infty}]=O\!\bigl((nm)^{-1/2}\bigr)$, the average exploitability obeys
$\frac{1}{T}\sum_{t=1}^{T}\!\text{Exploit}(\sigma_{t})\;\leq$ $\underbrace{O\!\Bigl(\tfrac{1}{T}\sqrt{\ln k_{T}\sum_{t=1}^{T}\!\|v_{t}-v_{t-1}\|_{\infty}^{2}}\Bigr)}_{\text{OMWU Regret}}\;+\;\underbrace{O\!\Bigl(\tfrac{1}{T}\sum_{t=1}^{T}(nm)^{-1/2}\Bigr)}_{\text{MC Estimation Error}}$ (10)
$\;+\;\underbrace{\tfrac{1}{T}\sum_{t=1}^{T}\varepsilon_{\mathrm{BR}}}_{\text{Amortized BR Error}}\;+\;\underbrace{O\!\Bigl(\tfrac{1}{T}\sum_{z\neq z^{\star}}\tfrac{\ln T}{\Delta_{z}}\Bigr)}_{\text{Oracle Regret}}.$
This bound corroborates the intuition behind Gems. As the simulation budget grows ($nm\!\to\!\infty$) and generator training improves ($\varepsilon_{\mathrm{BR}}\!\to\!0$), exploitability is ultimately driven by the no-regret property of the OMWU solver, achieving convergence guarantees competitive with traditional methods while avoiding quadratic complexity.
@@ 3.7 Generalization to $n$-Player and General-Sum Games
Although the Gems framework is introduced in the two-player zero-sum (2P-ZS) setting for clarity, the same principles extend naturally to the $n$-player general-sum (NP-GS) case. The amortized generator, Monte-Carlo rollouts, and bandit oracle remain intact; only the meta-game formulation generalizes.
@@
Full derivations, algorithmic details, and convergence proofs for the $n$-player extension appear in Appendix Part II.
@@ 3.8 Step-Size (ETA) Scheduler for the OMWU Meta-Update
$\eta_{t}\;=\;\begin{cases}\eta_{0},&\texttt{const},\\[4.0pt] \displaystyle\frac{\eta_{0}}{\sqrt{t}},&\texttt{sqrt},\\[10.0pt] \displaystyle\frac{\eta_{0}}{1+\alpha t},\ \alpha>0,&\texttt{harmonic},\end{cases}\qquad\text{with default }\eta_{0}=0.08,\ \alpha=0.5.$ (12)
@@
const keeps a fixed step size and adapts quickly to changes in the restricted game, but can overreact to noisy payoff estimates. sqrt decays gently ($\eta_{t}\propto t^{-1/2}$), a standard choice in online learning that balances responsiveness and variance. harmonic decays as $\eta_{t}\propto t^{-1}$, yielding more conservative late-stage updates and improved noise suppression; the reference implementation enables this schedule automatically when the slow preset is active (together with EMA smoothing and a period-gated oracle).
@@
Variation-aware regret with scheduled $\eta_{t}$.
@@
where $k_{T}$ is the (growing) number of anchors at time $T$. Equation 14 formalizes the tradeoff encoded by equation 12: larger $\eta_{t}$ reduces the stability term but amplifies sensitivity to payoff variation and sampling noise, while smaller $\eta_{t}$ does the reverse. In practice, sqrt performs well when the meta-game evolves moderately (frequent oracle additions), whereas harmonic is preferred in low-noise, late-phase training.
@@ 4.2 Equilibrium Finding in Kuhn Poker
We benchmark GEMS in Kuhn Poker [5], a classic imperfect information game where agents must learn mixed strategies involving bluffing. We evaluate performance using exploitability, which measures how close a policy is to the Nash Equilibrium (lower is better). GEMS is compared against a suite of strong PSRO variants over 40 training iterations for 5 seeds. This experiment is designed to assess GEMS’s ability to find near-optimal policies in an extensive-form game with imperfect information. The hypothesis is that GEMS’s architectural approach, which is exploring a continuous latent space with a single generator, will allow it to find low-exploitability strategies more efficiently and in fewer iterations than methods based on expanding a discrete set of policies.
Figure 3: Equilibrium Finding in Kuhn Poker. GEMS rapidly converges to a significantly lower exploitability than strong PSRO baselines (Left), while demonstrating superior efficiency in cumulative training time (Right).
@@
The results in Fig.3 show the convergence of exploitability for each algorithm. GEMS demonstrates the fastest and most direct convergence to a low-exploitability policy. By iteration 40, GEMS achieves an exploitability of approximately 0.159, significantly outperforming the next-best baseline, Double Oracle, which only reached an exploitability of 0.778. The other PSRO variants, while also showing improvement, converged at a considerably slower rate and achieved higher final exploitability within the 40-iteration budget. The superior performance in Kuhn Poker highlights GEMS’s strength in solving games that require nuanced, mixed strategies. The core of Kuhn Poker involves probabilistic actions like bluffing, which are difficult to represent as a simple combination of a few deterministic policies. We argue that GEMS’s single amortized generator, operating over a continuous latent space, is naturally suited to representing these complex mixed strategies. In contrast, methods that expand a discrete set of policies, like the PSRO family, must approximate a mixed strategy through a convex combination of many individual policies, which can require far more iterations to converge. The EB-UCB oracle effectively guides the search through the generator’s latent space to find strategically potent policies quickly. This experiment demonstrates that GEMS’s advantages extend beyond pure scalability to superior sample efficiency in solving canonical game-theoretic benchmarks.
@@ 4.3 Performance and Scalability on Multi-Agent Tag
We conduct our analysis in the simple_tag environment from PettingZoo [22], where three cooperative pursuers must learn coordinated strategies, such as flanking, to capture a faster evader. This benchmark is designed to reward sophisticated coordination while punishing naive "herding" policies. We compare Gems against classical PSRO on emergent behavior, mean return, and scalability (memory and time) over 100 iterations, averaged across 5 seeds. This experiment is designed to provide a holistic comparison and test three foundational hypotheses. First, that Gems overcomes the critical scalability bottlenecks of Psro. Second, that this efficiency does not come at a performance cost. Third, that Gems learns policies of a higher strategic quality, both quantitatively (as measured by mean return) and qualitatively (as observed in emergent agent coordination).
@@
Figure 4: Emergent Agent Trajectories in the Multi-Agent Tag Environment. Top row : GEMS, Bottom row : PSRO. This figure qualitatively compares the strategies learned by GEMS and classical PSRO. The top row shows that adversaries (red circles) trained with GEMS learn sophisticated, coordinated strategies like flanking and cornering to effectively trap the evader (green dot). In contrast, the bottom row shows that PSRO-trained agents adopt a less effective "herding" behavior, pursuing the target in a single, uncoordinated group. This clear difference in strategic complexity explains the superior performance and higher returns achieved by GEMS, as reflected in the quantitative results.
Figure 5: Performance and Scalability of GEMS vs. PSRO in Multi-Agent Tag. Compared to classical PSRO, GEMS achieves a higher and more stable mean return (Left), while maintaining a constant memory footprint (Middle) and near-linear cumulative training time (Right). The results show GEMS overcomes the core memory and computational bottlenecks of PSRO while learning more effective policies
4.4 More Environments
We refer the reader to Appendix Part III for more experiments.

### v2

@@ 3.5 Training the Generator with Amortized Best Response
Once a promising latent code $z_{t}^{\star}$ has been identified, Gems must ensure that the generator $G_{\theta}$ can realize the associated high-performing policy while retaining its ability to produce previously effective policies. This is achieved via an Amortized Best-Response with a Trust Region (ABR-TR) objective, inspired by trust-region methods in deep reinforcement learning such as TRPO and PPO (Schulman et al. (2015; 2017)).
Rather than training a new network from scratch, Gems fine-tunes the existing generator $G_{\theta}$ to maximize the performance of a curated set of promising latent codes, drawn from a distribution $q_{t}$, against the opponent mixture $\sigma_{t+1}$. A KL-divergence penalty against a frozen, older generator $\theta^{-}$ serves as a trust region, mitigating catastrophic forgetting:
$\mathcal{L}_{\mathrm{ABR-TR}}(\theta)\;=\;\mathbb{E}_{z\sim q_{t},\,j\sim\sigma_{t+1}}\Bigl[\widehat{A}_{\pi_{G_{\theta}(z)}}\;-\;\beta\,\operatorname{KL}\bigl(\pi_{G_{\theta}(z)}\;\|\;\pi_{G_{\theta^{-}}}(z)\bigr)\;-\;\lambda_{J}\|JG_{\theta}(z)\|_{F}^{2}\Bigr],$ (11)
where $\widehat{A}$ denotes a suitable advantage estimator, $\beta>0$ is a hyperparameter controlling trust-region strength and $JG_{\theta}(z)$ denotes the Jacobian matrix of the generator with respect to its input $z$. Taking a few gradient-ascent steps on $\mathcal{L}_{\mathrm{ABR-TR}}$ each iteration amortizes best-response computation while preserving a stable, single generator.
@@ 3.6 Overall Exploitability Bound
The preceding components now combine to yield a convergence guarantee for the entire Gems algorithm. To make this composition explicit, we first state a decomposition result that separates the average exploitability into four interpretable sources of error:
  • inherent regret of the OMWU meta-strategy solver (under time-varying payoffs),
  • noise from Monte Carlo payoff estimation,
  • sub-optimality from bandit-based oracle (anchor) selection,
  • approximation error induced by amortized generator training.
Amortized best-response gap.
Let $\pi_{\mathrm{BR}}(\pi_{j})$ be the true best response to an opponent policy $\pi_{j}$. We define the amortized best-response error as the worst-case expected gap (over iterations) between a true best response and the policy produced by the generator at the selected latent code $z_{t}^{\star}$:
$\varepsilon_{\mathrm{BR}}\;=\;\sup_{t}\mathbb{E}_{j\sim\sigma_{t}}\Bigl[r\bigl(\pi_{\mathrm{BR}}(j),\pi_{j}\bigr)\;-\;r\bigl(\pi_{G_{\theta}(z_{t}^{\star})},\pi_{j}\bigr)\Bigr]\;\geq\;0.$ (12)
Exploitability decomposition.
Let $M$ denote the (implicit) meta-game payoff operator and define the true payoff vector $v_{t}:=M\sigma_{t}$ at iteration $t$. Let $\hat{v}_{t}$ be the Monte-Carlo estimator of $v_{t}$ computed from $nm$ rollouts per queried matchup (as in Section 3.2). Let $z_{t}^{\star}$ be the (possibly approximate) anchor selected by the EB-UCB oracle at iteration $t$, and let $\Delta_{z}$ denote the suboptimality gap for anchor $z$.
Proposition 3.4 (Exploitability decomposition).
Assume rewards lie in $[0,1]$. At each iteration $t$, the instantaneous exploitability satisfies
$\text{Exploit}(\sigma_{t})\;\leq\;\underbrace{\bigl\langle v_{t}-v_{t-1},\,\sigma_{t}-\sigma_{t}^{\star}\bigr\rangle}_{\text{OMWU meta-dynamics term}}\;+\;\underbrace{2\|\hat{v}_{t}-v_{t}\|_{\infty}}_{\text{MC estimation error}}\;+\;\underbrace{\varepsilon_{\mathrm{BR}}}_{\text{amortized BR error}}\;+\;\underbrace{\Bigl(\max_{z}v_{t}(z)-v_{t}(z_{t}^{\star})\Bigr)}_{\text{oracle suboptimality}}.$ (13)
Here $\sigma_{t}^{\star}$ denotes the best fixed meta-strategy in hindsight for the (time-varying) payoff sequence.
Proof sketch. Starting from the definition of exploitability as a best-response gap against $\sigma_{t}$, we add and subtract (i) Monte-Carlo payoffs $\hat{v}_{t}$, (ii) the best-anchor payoff $\max_{z}v_{t}(z)$, and (iii) the generator-induced best response at $z_{t}^{\star}$. Triangle inequalities yield the MC term and the amortized BR term, while the oracle term captures the loss from selecting $z_{t}^{\star}$ instead of $\arg\max_{z}v_{t}(z)$. The remaining meta-dynamics term is controlled by the OMWU guarantee for time-varying payoffs.
Overall finite-population bound.
We now upper bound each term in Proposition 3.4 using the corresponding component guarantees.
@@
Assume rewards lie in $[0,1]$ and that oracle selection operates on a two-time-scale schedule. With OMWU step size $\eta=\Theta\bigl(\sqrt{\ln k_{T}/T}\bigr)$ and Monte-Carlo budget satisfying $\mathbb{E}[\|\hat{v}_{t}-v_{t}\|_{\infty}]=O\bigl((nm)^{-1/2}\bigr)$ for each $t$, the average exploitability obeys
$\frac{1}{T}\sum_{t=1}^{T}\text{Exploit}(\sigma_{t})\;\leq$ $\underbrace{O\Bigl(\tfrac{1}{T}\sqrt{\ln k_{T}\sum_{t=1}^{T}\|v_{t}-v_{t-1}\|_{\infty}^{2}}\Bigr)}_{\text{OMWU regret}}\;+\;\underbrace{O\bigl((nm)^{-1/2}\bigr)}_{\text{MC estimation error}}$ (14)
$\;+\;\underbrace{\varepsilon_{\mathrm{BR}}}_{\text{amortized BR error}}\;+\;\underbrace{O\Bigl(\tfrac{1}{T}\sum_{z\neq z^{\star}}\tfrac{\ln(T/\delta)}{\Delta_{z}}\Bigr)}_{\text{oracle regret}}.$
Proof (one line). Apply Proposition 3.4, then substitute the bounds from the OMWU time-varying regret guarantee, the Monte-Carlo concentration bound for $\|\hat{v}_{t}-v_{t}\|_{\infty}$, and the EB-UCB oracle regret bound, and average over $t=1,\dots,T$.
This bound corroborates the intuition behind Gems. As the simulation budget grows ($nm\to\infty$) and generator training improves ($\varepsilon_{\mathrm{BR}}\to 0$), exploitability is ultimately driven by the no-regret property of the OMWU solver, achieving convergence guarantees competitive with traditional methods while avoiding explicit quadratic meta-game storage.
@@ 3.7 Generalization to n-Player and General-Sum Games
Although the Gems framework is introduced in the two-player zero-sum (2P-ZS) setting for clarity, the same principles extend naturally to the $n$-player general-sum (NP-GS) case. The amortized generator, Monte Carlo rollouts, and bandit oracle remain intact; only the meta-game formulation generalizes.
@@
Full derivations, algorithmic details, and convergence proofs for the $n$-player extension appear in Appendix II.
@@ 3.8 Step-Size (ETA) Scheduler for the OMWU Meta-Update
$\eta_{t}\;=\;\begin{cases}\eta_{0},&\text{const},\\[4.0pt] \displaystyle\frac{\eta_{0}}{\sqrt{t}},&\text{sqrt},\\[10.0pt] \displaystyle\frac{\eta_{0}}{1+\alpha t},\ \alpha>0,&\text{harmonic},\end{cases}\qquad\text{with default }\eta_{0}=0.08,\ \alpha=0.5.$ (16)
@@
❶‘const’ keeps a fixed step size and adapts quickly to changes in the restricted game, but can overreact to noisy payoff estimates. ❷ ‘sqrt’ decays gently ($\eta_{t}\propto t^{-1/2}$), a standard choice in online learning that balances responsiveness and variance. ❸ ‘harmonic’ decays as $\eta_{t}\propto t^{-1}$, yielding more conservative late-stage updates and improved noise suppression.
Require :  Initial anchor set $\mathcal{Z}_{0}=\{z_{1},\dots,z_{k_{0}}\}$, Generator $G_{\theta}$, Meta-strategy $\sigma_{0}$ (uniform), Horizon $T$, Budgets $N_{MC},B,M$, Steps $K_{ABR}$, Rates $\eta,\alpha$, Coeffs $\beta,\lambda_{J}$.
Output : Trained Generator $G_{\theta}$, Final Meta-strategy $\sigma_{T}$, Anchor set $\mathcal{Z}_{T}$.
1ex
for $t=1,\dots,T$ do
 # Phase 1: Meta-Game Estimation (Monte Carlo)
   Sample opponent batch $\{j_{s}\}_{s=1}^{N_{MC}}\sim\sigma_{t-1}$
   Compute empirical payoff vector $\hat{v}_{t}\in\mathbb{R}^{|\mathcal{Z}_{t-1}|}$ via Eq. (2):
 $\hat{v}_{t,i}\leftarrow\frac{1}{N_{MC}m}\sum_{s=1}^{N_{MC}}\sum_{l=1}^{m}\mathcal{R}(\pi_{G_{\theta}(z_{i})},\pi_{G_{\theta}(z_{j_{s}})})$
  Estimate mean game value $\hat{\overline{r}}_{t}$ via $B$ joint samples $(i,j)\sim\sigma_{t-1}\times\sigma_{t-1}$
  1ex
 # Phase 2: Meta-Strategy Update (Optimistic MWU)
   Compute optimistic gradient estimate (with $\hat{v}_{0}=\mathbf{0}$):
 $g_{t}\leftarrow 2\hat{v}_{t}-\hat{v}_{t-1}-\hat{\overline{r}}_{t}\mathbf{1}$
   Update meta-strategy probabilities $\forall i\in\{1,\dots,|\mathcal{Z}_{t-1}|\}$:
 $\sigma_{t}(i)\propto\sigma_{t-1}(i)\cdot\exp\left(\eta\cdot g_{t}(i)\right)$
  1ex
 # Phase 3: Population Expansion (EB-UCB Oracle)
   Generate candidate pool $\Lambda_{t}$ via stochastic mutation of top anchors
 for each candidate $z\in\Lambda_{t}$ do
      Estimate mean return $\hat{\mu}_{t}(z)$ and variance $\widehat{\mathrm{Var}}_{t}(z)$ vs $\sigma_{t}$
      Calculate acquisition score $\mathrm{UCB}(z)$ via Eq. (7):
    $\mathrm{UCB}(z)\leftarrow\hat{\mu}_{t}(z)+\sqrt{\frac{2\widehat{\mathrm{Var}}_{t}(z)\ln(3/\delta)}{M}}-\lambda_{J}\|JG_{\theta}(z)\|_{F}^{2}$
  Select optimal candidate: $z^{*}_{t}\leftarrow\arg\max_{z\in\Lambda_{t}}\mathrm{UCB}(z)$
   Update anchor set: $\mathcal{Z}_{t}\leftarrow\mathcal{Z}_{t-1}\cup\{z^{*}_{t}\}$
  1ex
 # Phase 4: Amortized Best-Response (ABR-TR)
   Store current generator state: $\theta_{\text{old}}\leftarrow\theta$
 for $k=1,\dots,K_{ABR}$ do
      Sample batch of anchors $z\sim\mathcal{Z}_{t}$ and opponents $j\sim\sigma_{t}$
      Compute Advantages $\hat{A}_{\pi_{G_{\theta}(z)}}$ via GAE
      Update $\theta$ by ascending on regularized objective (Eq. 11):
    $\theta\leftarrow\theta+\alpha\nabla_{\theta}\mathbb{E}\Big[\hat{A}-\beta D_{KL}(\pi_{G_{\theta}}||\pi_{G_{\theta_{\text{old}}}})-\lambda_{J}\|JG_{\theta}(z)\|_{F}^{2}\Big]$
return $G_{\theta},\sigma_{T},\mathcal{Z}_{T}$
Algorithm 1 Generative Evolutionary Meta-Solver (Gems)
@@
Variation-aware regret with scheduled ($\eta_{t}$).
@@
where $k_{T}$ is the (growing) number of anchors at time $T$. Equation 18 formalizes the tradeoff encoded by equation 16: larger $\eta_{t}$ reduces the stability term but amplifies sensitivity to payoff variation and sampling noise, while smaller $\eta_{t}$ does the reverse. In practice, ‘sqrt’ performs well when the meta-game evolves moderately (frequent oracle additions), whereas ‘harmonic’ is preferred in low-noise, late-phase training.
@@
3.9 Algorithm
Algorithm 1 provides the full procedural realization of the method described in §3. Each iteration consists of four phases that mirror the preceding subsections: (i) Monte Carlo estimation of the meta-game values (Eq. equation 2), (ii) an OMWU meta-update using the optimistic hint (Eq. equation 17), (iii) population expansion via EB-UCB anchor selection (Eq. (7) and Eq. equation 13), and (iv) amortized generator training with a KL trust region (Eq. equation 11). The only persistent state is the generator parameters $\theta$, the anchor set $\mathcal{Z}_{t}$, and the meta-strategy $\sigma_{t}$, enabling Gems to refine equilibria while avoiding explicit policy enumeration or payoff-matrix storage.
@@ 4.2 Equilibrium Finding in Kuhn Poker
We benchmark Gems in Kuhn Poker (Kuhn, 2016), a classic imperfect information game where agents must learn mixed strategies involving bluffing. We evaluate performance using exploitability, which measures how close a policy is to the Nash Equilibrium (lower is better). Gems is compared against a suite of strong Psro variants over 40 training iterations for 5 seeds. This experiment is designed to assess Gems’s ability to find near-optimal policies in an extensive-form game with imperfect information. The hypothesis is that Gems’s architectural approach, which is exploring a continuous latent space with a single generator, will allow it to find low-exploitability strategies more efficiently and in fewer iterations than methods based on expanding a discrete set of policies.
Figure 4: Equilibrium Finding in Kuhn Poker over 5 seeds [0–4]. Gems rapidly converges to a significantly lower exploitability than strong Psro baselines and NeuPL (Left), while demonstrating efficiency in cumulative training time (Right).
@@
The results in Fig. 4 show the convergence of exploitability for each algorithm. Gems demonstrates the fastest and most direct convergence to a low-exploitability policy. By iteration 40, Gems achieves an exploitability of approximately 0.18, significantly outperforming the next-best baseline, E-psro, which only reached an exploitability of 0.44. The other Psro variants, while also showing improvement, converged at a considerably slower rate and achieved higher final exploitability within the 40-iteration budget. The superior performance in Kuhn Poker highlights Gems’s strength in solving games that require nuanced, mixed strategies. The core of Kuhn Poker involves probabilistic actions like bluffing, which are difficult to represent as a simple combination of a few deterministic policies. We argue that Gems’s single amortized generator, operating over a continuous latent space, is naturally suited to representing these complex mixed strategies. In contrast, methods that expand a discrete set of policies, like the Psro family, must approximate a mixed strategy through a convex combination of many individual policies, which can require far more iterations to converge. The EB-UCB oracle effectively guides the search through the generator’s latent space to find strategically potent policies quickly. This experiment demonstrates that Gems’s advantages extend beyond pure scalability to superior sample efficiency in solving canonical game-theoretic benchmarks.
@@ 4.3 Performance and Scalability on Multi-Agent Tag
We conduct our analysis in the Simple Tag environment from PettingZoo (Terry et al., 2021), where three cooperative pursuers must learn coordinated strategies, such as flanking, to capture a faster evader. This benchmark is designed to reward sophisticated coordination while punishing naive “herding” policies. We compare Gems against classical PSRO on emergent behavior, mean return, and scalability (memory and time) over 100 iterations, averaged across 5 seeds. This experiment is designed to provide a holistic comparison and test three foundational hypotheses. First, that Gems overcomes the critical scalability bottlenecks of Psro. Second, that this efficiency does not come at a performance cost. Third, that Gems learns policies of a higher strategic quality, both quantitatively (as measured by mean return) and qualitatively (as observed in emergent agent coordination).
@@
GEMS
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
PSRO
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 5: Emergent agent trajectories in the multi-agent tag environment. Top row: Gems. Bottom row: Psro. Columns show uniformly sampled frames from a single rollout (frames 0, 10, 20, 30, 40, 50 of the 50-frame GIF). This figure qualitatively compares the strategies learned by GEMS and classical PSRO. The top row shows that adversaries (red circles) trained with GEMS learn sophisticated, coordinated strategies like flanking and cornering to effectively trap the evader (green dot). In contrast, the bottom row shows that PSRO-trained agents adopt a less effective "herding" behavior, pursuing the target in a single, uncoordinated group. This clear difference in strategic complexity is consistent with the superior performance and higher returns achieved by GEMS, as reflected in the quantitative results.
Figure 6: Performance and Scalability of Gems vs. Psro in Multi-Agent Tag. Compared to classical Psro, Gems achieves a higher and more stable mean return (Left), while maintaining a constant memory footprint (Middle) and near-linear cumulative training time (Right). The results show Gems overcomes the core memory and computational bottlenecks of Gems while learning more effective policies